The square of an integer is 66 more then five times the integer. Find two numbers for which this is true.
Answer by ankor@dixie-net.com(22740)

The square of an integer is 66 more then five times the integer.
Find two numbers for which this is true.
:
x^2 - 5x = 66
A quadratic equation
x^2 - 5x - 66 = 0
Factors to
(x-11)(x+6) = 0
Two solutions
x=11
and
x=-6
:
Check x=-6 in original equation
-6^2 - 5(-6) =
36 - (-30) =
36 + 30 = 66
:
You can check x=11